Prove that (1 * 2) + (2 * 3) + (3 * 4) + (4 * 5) + ....+ n (n + 1) = n (n + 1) (n + 2) / 3 for n positive integers
1
Expert's answer
2020-12-17T19:16:56-0500

"\text{We use the method of mathematical induction}"

"\text{for }n=1"

"n(n+1)= 1*2 =2"

"\frac{n(n+1)(n+2)}{3}=\frac{1*2*3}{3}=2"

"\text{for }n=1 \text{ statement is true}"

"\text{suppose the statement is true for } n =k"

"(1 * 2) + (2 * 3) + ....+ k (k + 1) = \frac{k (k + 1) (k + 2) }{3}\ (1)"


"\text{сonsider the expression for } n = k+1"

"(1 * 2) + (2 * 3) + ....+ k (k + 1)+(k+1)(k+2)" =

"\text{from equality (1):}"

"=\frac{k (k + 1) (k + 2) }{3}+(k+1)(k+2)= \frac{k (k + 1) (k + 2)+3(k+1)(k+2) }{3}=\frac{ (k + 1) (k + 2)(k+3) }{3}"

"\text{i.e}"

"(1 * 2) + (2 * 3) + ....+ k (k + 1)+(k+1)(k+2)=" "\frac{ (k + 1) (k + 2)(k+3) }{3}"

"\text{ the statement is true for } n =k+1"


"\text{the expression is true for n = 1 on the assumption}"

"\text{that the expression is true for n = k follows that}"

"\text{the expression is also true for n= k+1}"

"\text{therefore, the expression is true for all positive numbers}"


Answer:the statement under the conditions of the problem is correct,

proven by mathematical induction